Calculate the weight percent of ascorbic acid in a tablet of Vitamin C from the following data: A 100 mg sample of a crushed Vitamin C tablet was dissolved in 40 mL of H2SO4 and 20 mL of water. Two grams of KI and 35 mL of 0.0107 M KIO3 solution was added, and the mixture titrated to a starch endpoint. The titration required 12 mL of 0.0790 M thiosulfate solution.
Answer #1

The no. of moles of KIO3 = (35/1000) L * 0.0107 mol/L = 0.0003745 mol

KIO3 + 5KI + 6H+ ----> 3I2 + 3H2O + 6K+

Therefore, the no. of moles of I2 = 3*0.0003745 = 0.0011235 mol

The no. of molesof Na2S2O3 = (12/1000) L * 0.079 mol/L = 0.000948 mol

2Na2S2O3 + I2 ----> Na2S4O6 + 2I-

The no. of moles of I2 consumed in the above reaction = 0.000948/2 = 0.000474 mol

Now, the no. of moles of I2 remaining = 0.0011235 - 0.000474 = 0.0006495 mol

C6H8O6 + I2 ----> C6H6O6 + 2I- + 2H+

Therefore, the no. of moles of ascorbic acid in the vitamin C tablet = 0.000474 mol

i.e. The mass of ascorbic acid in the vitamin C tablet = 0.000474 mol * 176 g/mol = 0.083 g or 83 mg

Hence, the weight percent of ascorbic acid in a tablet of vitamin C = 83%